# Zhang, Liu, and Zhao

### v3

@@ 9 Surrogation Construction of $F(\cdot;\xi)$ for Case 1 to Case 3
$\widehat{\kappa}\coloneqq\limsup_{\theta\rightarrow\theta^{\prime},\theta\neq\theta^{\prime}}\frac{2(\widehat{F}^{\varepsilon}_{\min}(\theta;\theta^{\prime},\xi)-F(\theta;\xi))}{\left\lVert\theta-\theta^{\prime}\right\rVert^{2}}<+\infty.$
For the piecewise affine function $f(\theta;\xi)\coloneqq g(\theta;\xi)-h(\theta;\xi)$, we have $\mathcal{I}_{h}^{0}(\theta)\subseteq\mathcal{I}_{h}^{\varepsilon}(\theta^{\prime})$ when $\theta\in\Theta$ is sufficiently close to $\theta^{\prime}$. It follows from (P2) that for any $\theta\in\Theta$ sufficiently close to $\theta^{\prime}$,
$f(\theta;\xi)\leq\min_{I_{2}\in\mathcal{I}^{\varepsilon}_{h}(\theta^{\prime})}\widehat{f}(\theta;\xi,I_{2})\leq\min_{I_{2}\in\mathcal{I}_{h}^{0}(\theta)}\widehat{f}(\theta;\xi,I_{2})=f(\theta;\xi).$
Therefore, by similar arguments for lower surrogation functions of $f(\cdot;\xi)$, we obtain that for any $\theta\in\Theta$ sufficiently close to $\theta^{\prime}$,
$\min_{I_{2}\in\mathcal{I}^{\varepsilon}_{h}(\theta^{\prime})}\widehat{f}(\theta;\xi,I_{2})=f(\theta;\xi)=\max_{I_{1}\in\mathcal{I}^{\varepsilon}_{g}(\theta^{\prime})}\widecheck{f}(\theta;\xi,I_{1}).$
For Case 1, since the objective is piecewise affine, we directly obtain $\widehat{\kappa}=0$. For Case 2, when $\theta$ is sufficiently close to $\theta^{\prime}$, we have
$\widehat{F}^{\varepsilon}_{\min}(\theta;\theta^{\prime},\xi)$ $\coloneqq\min_{(I_{1},I_{2})\in\mathcal{I}^{\varepsilon}(\theta^{\prime})}\widehat{F}(\theta;\theta^{\prime},\xi,(I_{1},I_{2}))$
$=\min_{I_{2}\in\mathcal{I}^{\varepsilon}_{h}(\theta^{\prime})}\varphi^{\uparrow}(\widehat{f}(\theta;\xi,I_{2});\xi)+\min_{I_{1}\in\mathcal{I}^{\varepsilon}_{g}(\theta^{\prime})}\varphi^{\downarrow}(\widecheck{f}(\theta;\xi,I_{1});\xi)$
$=\varphi^{\uparrow}\left(\min_{I_{2}\in\mathcal{I}^{\varepsilon}_{h}(\theta^{\prime})}\widehat{f}(\theta;\xi,I_{2});\xi\right)+\varphi^{\downarrow}\left(\max_{I_{1}\in\mathcal{I}^{\varepsilon}_{g}(\theta^{\prime})}\widecheck{f}(\theta;\xi,I_{1});\xi\right)$
$=F(\theta;\xi),$
also implying $\widehat{\kappa}=0$. For Case 3, when $\theta$ is sufficiently close to $\theta^{\prime}$, we obtain
@@
$\widehat{F}^{\varepsilon}_{\min}(\theta;\theta^{\prime},\xi)=\nabla\varphi(f(\theta^{\prime};\xi);\xi)^{\top}(f(\theta;\xi)-f(\theta^{\prime};\xi))+\varphi(f(\theta^{\prime};\xi);\xi)+\frac{1}{2}L_{\nabla\varphi}L_{f}^{2}\left\lVert\theta-\theta^{\prime}\right\rVert^{2}.$
Thus, by the second-order Taylor expansion of $\varphi(\cdot;\xi)$ at $f(\theta^{\prime};\xi)$, we obtain
@@
$\widehat{F}^{\varepsilon}_{\min}(\theta;\theta^{\prime},\xi)-F(\theta;\xi)=\frac{1}{2}L_{\nabla\varphi}L_{f}^{2}\left\lVert\theta-\theta^{\prime}\right\rVert^{2}-\frac{1}{2}v^{\top}\nabla^{2}\varphi(f(\theta^{\prime};\xi);\xi)v+o(\left\lVert v\right\rVert^{2}),$
@@
10 Details of Convergence Analysis in Section 4
@@
Proof 10.1
Proof of Proposition 4.2. Given $\varepsilon\geq 0$ and $\rho>0$, for any $\theta^{*}\in\mathcal{S}^{*}$ it holds that $\widehat{F}^{\varepsilon}_{\min}(\theta;\theta^{*})+\frac{\rho}{2}\left\lVert\theta-\theta^{*}\right\rVert^{2}>F(\theta)\geq F(\theta^{*})=\widehat{F}^{\varepsilon}_{\min}(\theta^{*};\theta^{*})$ for any $\theta\neq\theta^{*}$, showing $\theta^{*}\in\mathcal{S}^{d}_{\varepsilon,\rho}$ and thus $\mathcal{S}^{*}\subseteq\mathcal{S}^{d}_{\varepsilon,\rho}$. For any $\theta^{d}\in\mathcal{S}^{d}_{\varepsilon,\rho}$, from Condition (P4) and the first-order optimality condition, it follows that $F^{\prime}(\theta^{d};\theta-\theta^{d})=\widehat{F}^{\varepsilon}_{\min}(\cdot;\theta^{d})^{\prime}(\theta^{d};\theta-\theta^{d})\geq 0$ for all $\theta\in\Theta$, yielding $\theta^{d}\in\mathcal{S}^{d}$ and thus $\mathcal{S}^{d}_{\varepsilon,\rho}\subseteq\mathcal{S}^{d}$.
It remains to show $\mathcal{S}^{d}\subseteq\mathcal{S}^{d}_{0,\rho}$ for any $\rho>0$. Assume by contradiction that there exists $\theta^{d}\in\mathcal{S}^{d}\backslash\mathcal{S}^{d}_{0,\widehat{\rho}}$ for some $\widehat{\rho}>0$. Then there exist $\widehat{I}\in\mathcal{I}^{0}(\theta^{d})$ and $\widehat{\theta}(\neq\theta^{d})\in\Theta$ such that $\widehat{\theta}\in\mathop{\arg\min}_{\theta\in\Theta}\left\{\widehat{F}(\theta;\theta^{d},\widehat{I}\,)+\frac{\widehat{\rho}}{2}\left\lVert\theta-\theta^{d}\right\rVert^{2}\right\}$. By the strong convexity of the function $\theta\mapsto\widehat{F}(\theta;\theta^{d},\widehat{I}\,)+\frac{\widehat{\rho}}{2}\,\left\lVert\theta-\theta^{d}\right\rVert^{2}$, we have $\widehat{F}(\cdot;\theta^{d},\widehat{I})^{\prime}(\theta^{d};\widehat{\theta}-\theta^{d})<0$. Thus, there exists $t_{0}>0$ such that
$F(\theta^{d}+t(\widehat{\theta}-\theta^{d}))\leq\widehat{F}(\theta^{d}+t(\widehat{\theta}-\theta^{d});\theta^{d},\widehat{I})<\widehat{F}(\theta^{d};\theta^{d},\widehat{I})=F(\theta^{d}),\quad\forall\,t\in(0,t_{0}],$
@@
The assumption is naturally satisfied when the outer function of $F$ is in Case 1 or Case 2. Specifically, from the proof of Claim 2 afterward, we see that in these two cases, $\Theta$ can be decomposed into finitely many convex regions, on each of which any d-stationary point $\theta^{d}\in\mathcal{S}^{d}$ is a minimizer of $F$ over that region. Since the number of such regions is finite, the function $F$ takes finitely many values on $\mathcal{S}^{d}$, verifying Assumption 1. Discussions on the finite number of d-stationary values for more general nonconvex problems with piecewise and composite structures can be seen in Cui and Pang (2018).
@@
For problem (9) with $\max_{\theta\in\Theta}\left\lVert\theta\right\rVert_{\infty}=\mu$ and Assumption 1 holding for the surrogate family $\{\widehat{F}(\cdot;\theta^{\prime},\xi,I)\}_{I\in\mathcal{I}^{\varepsilon}(\theta^{\prime})}$ of $F(\cdot;\xi)$ for each $\xi\in S$, the following statements hold:
@@
    With $\varphi$ in Case 1 or Case 2, $\mathcal{S}^{d}_{\varepsilon,\rho}=\mathcal{S}^{*}$ for any $\varepsilon\in[\varepsilon_{0},+\infty)$ and $\rho\in(0,\frac{M_{\Delta}}{2q\mu^{2}}]$, where $\varepsilon_{0}\coloneqq 2\mu\sqrt{q}\sqrt{\max_{s\in[n]}\left\lVert x^{s}\right\rVert^{2}+1}$, and $M_{\Delta}$ is the constant defined in 1.

### v4

@@ 7 Surrogation Construction of $F(\cdot,\xi)$ for Case 1 to Case 3
    Quadratic surrogation gap: for some $\kappa>0$, $F(\theta,\xi)+\displaystyle{\frac{\kappa}{2}}\left\lVert\theta-\bar{\theta}\right\rVert^{2}\geq\widehat{F}^{\varepsilon}_{\min}(\theta,\xi;\bar{\theta})$ for any $\theta\in\Theta$.
For directionally differentiable surrogate functions, Condition (P4) is stronger than directional derivative consistency condition because it also limits the growth rate of the surrogate function with respect to the original function, which is essential to develop the nonasymptotic convergence analysis in Section 8.4.
Proof 7.1
Proof of Proposition 3.2
We first construct the surrogate functions for the three cases and verify Assumptions (B1)-(B3), followed by the verification of Assumption (B4).
Case 1: Suppose that $\varphi(z,\xi)=\varphi_{1}(z,\xi)-\varphi_{2}(z,\xi)$, where $\varphi_{1}(z,\xi)=\max_{i\in[K_{\varphi}]}\varphi_{1i}(z,\xi)$ and $\varphi_{2}(z,\xi)=\max_{i\in[K_{\varphi}]}\varphi_{2i}(z,\xi)$ are both max-affine functions. We denote the affine functions by $\{\varphi_{1i}(z,\xi)=a_{1i}^{\top}z+b_{1i}\}_{i\in[K_{\varphi}]}$ and $\{\varphi_{2i}(z,\xi)=a_{2i}^{\top}z+b_{2i}\}_{i\in[K_{\varphi}]}$, where $\xi$ represents the random components of $\{a_{1i},b_{1i},a_{2i},b_{2i}\}_{i\in[K_{\varphi}]}$. Then the surrogate function of $F(\cdot,\xi)$ at $\bar{\theta}$ is constructed as follows,
$\widehat{F}(\theta,\xi;\bar{\theta},I)$ $=\max_{i\in[K_{\varphi}]}\left\{\max\{a_{1i},0\}^{\top}\widehat{f}(\theta,\xi;I_{2})-\max\{-a_{1i},0\}^{\top}\widecheck{f}(\theta,\xi;I_{1})+b_{1i}\right\}$
$\phantom{=}-\left(\max\{a_{2I_{3}},0\}^{\top}\widecheck{f}(\theta,\xi;I_{1})-\max\{-a_{2I_{3}},0\}^{\top}\widehat{f}(\theta,\xi;I_{2})+b_{2I_{3}}\right),$
where $I_{1}\in\mathcal{I}_{h}^{\varepsilon}(\bar{\theta},\xi)$, $I_{2}\in\mathcal{I}_{g}^{\varepsilon}(\bar{\theta},\xi)$, and $I_{3}\in\mathcal{I}_{\varphi_{2}}^{\varepsilon}(f(\bar{\theta},\xi),\xi)$. With a alight abuse of notation, we denote $I=(I_{1},I_{2},I_{3})\in\mathcal{I}^{\varepsilon}(\bar{\theta},\xi)\coloneqq\mathcal{I}_{g}^{\varepsilon}(\bar{\theta},\xi)\times\mathcal{I}_{h}^{\varepsilon}(\bar{\theta},\xi)\times\mathcal{I}_{\varphi_{2}}^{\varepsilon}(f(\bar{\theta},\xi),\xi)$. Assumptions (B1) and (B2) hold in this case by construction. Assumption (B3) is justified since given a fixed surrogation index $I$, the surrogate function is independent of the reference point $\bar{\theta}$.
Case 2: by assumption, $\varphi(\cdot,\xi)$ can be written as $\varphi(\cdot,\xi)=\varphi^{\uparrow}(\cdot,\xi)+\varphi^{\downarrow}(\cdot,\xi)$, where $\varphi^{\uparrow}(\cdot,\xi)$ and $\varphi^{\downarrow}(\cdot,\xi)$ are nondecreasing and nonincreasing convex functions respectively, that is, for any $\theta_{1},\theta_{2}\in\Theta$ satisfying $\theta_{1}\leq\theta_{2}$ componentwisely, we have $\varphi^{\uparrow}(\theta_{2},\xi)\geq\varphi^{\uparrow}(\theta_{1},\xi)$ and $\varphi^{\downarrow}(\theta_{2},\xi)\leq\varphi^{\downarrow}(\theta_{1},\xi)$. With $I=(I_{1},I_{2})\in\mathcal{I}^{\varepsilon}(\bar{\theta},\xi)$, the surrogate function of $F(\cdot,\xi)$ at $\bar{\theta}$ is constructed as
$\widehat{F}(\theta,\xi;\bar{\theta},I)=\varphi^{\uparrow}(\widehat{f}(\theta,\xi;I_{2}),\xi)+\varphi^{\downarrow}(\widecheck{f}(\theta,\xi;I_{1}),\xi),$
where $\widehat{f}(\theta,\xi;I_{2})$ and $\widecheck{f}(\theta,\xi;I_{1})$ are the upper and lower surrogate functions of $f(\theta,\xi)$ respectively. By monotonicity of $\varphi^{\uparrow}$ and $\varphi^{\downarrow}$, $\widehat{F}(\theta,\xi;\bar{\theta},I)\geq\varphi^{\uparrow}(f(\theta,\xi),\xi)+\varphi^{\downarrow}(f(\theta,\xi),\xi)=F(\theta,\xi)$ for any $\theta\in\Theta$, and thus the majorizing condition (P2) holds. Moreover, $\widehat{F}(\cdot,\xi;\bar{\theta},I)$ satisfies the convex condition (P3) since a nondecreasing convex function composite with a convex function and a nonincreasing convex function composite with a concave function are both convex. The touching condition (P1) clearly holds when $\varepsilon=0$. With a fixed index $I$, $\widehat{F}(\cdot,\xi;\bar{\theta},I)$ is also independent of $\bar{\theta}$ in this case, thus validating the bi-variate continuity assumption (B3).
Case 3: since $\Xi$ is finite, for any $\xi\in\Xi$, $\varphi(\cdot,\xi)$ is Lipschitz smooth with uniform modulus denoted by $L_{\nabla\varphi}$, and $f(\cdot,\xi)$ is Lipschitz continuous with uniform modulus $L_{f}\coloneqq 2\sqrt{\max_{s\in[n]}\left\lVert x^{s}\right\rVert^{2}+1}$. Then the surrogate function of $F(\cdot,\xi)$ at $\bar{\theta}$ with $I=(I_{1},I_{2})\in\mathcal{I}^{\varepsilon}(\bar{\theta},\xi)$ is constructed by surrogating the inner PA function according to componentwise signs of the gradient $\nabla\varphi(f(\bar{\theta},\xi),\xi)$ as follows,
$\widehat{F}(\theta,\xi;\bar{\theta},I)$ $=\max\{\nabla{\varphi}(f(\bar{\theta},\xi),\xi),0\}^{\top}\widehat{f}(\theta,\xi;I_{2})-\max\{-\nabla{\varphi}(f(\bar{\theta},\xi),\xi),0\}^{\top}\widecheck{f}(\theta,\xi;I_{1})$
@@
$\phantom{=,}-\nabla{\varphi}(f(\bar{\theta},\xi),\xi)^{\top}f(\bar{\theta},\xi)+\varphi(f(\bar{\theta},\xi),\xi)+\frac{1}{2}L_{\nabla\varphi}L_{f}^{2}\left\lVert\theta-\bar{\theta}\right\rVert^{2}.$
The convex condition (P3) holds evidently; the touching condition (P1) also holds when $\varepsilon=0$. The surrogate $\widehat{F}(\cdot,\xi;\cdot,I)$ is continuous on $\Theta\times\Theta$ since $\nabla\varphi(\cdot,\xi)$ is continuous, validating Assumption (B3). Based on the quadratic upper bound of a Lipschitz smooth function, the majorizing condition (P2) is verified as follows. Since $\varphi(\cdot,\xi)$ is $L_{\nabla\varphi}$-smooth, we have for any $z_{1},z_{2}\in\mathbb{R}^{d}$ that
$\varphi(z_{1},\xi)\leq\varphi(z_{2},\xi)+\nabla\varphi(z_{2},\xi)^{\top}(z_{1}-z_{2})+\frac{L_{\nabla\varphi}}{2}\left\lVert z_{1}-z_{2}\right\rVert^{2}.$
By setting $z_{1}=f(\theta,\xi)$ and $z_{2}=f(\bar{\theta},\xi)$, we obtain that for any $I=(I_{1},I_{2})\in\mathcal{I}^{\varepsilon}(\bar{\theta},\xi)$,
@@
$F(\theta,\xi)$ $\leq F(\bar{\theta},\xi)+\nabla\varphi(f(\bar{\theta},\xi),\xi)^{\top}(f(\theta,\xi)-f(\bar{\theta},\xi))+\frac{L_{\nabla\varphi}}{2}\left\lVert f(\theta,\xi)-f(\bar{\theta},\xi)\right\rVert^{2}$
$\leq F(\bar{\theta},\xi)+\max\{\nabla{\varphi}(f(\bar{\theta},\xi),\xi),0\}^{\top}\widehat{f}(\theta,\xi;I_{2})$
$\phantom{\leq,}-\max\{-\nabla{\varphi}(f(\bar{\theta},\xi),\xi),0\}^{\top}\widecheck{f}(\theta,\xi;I_{1})-\nabla\varphi(f(\bar{\theta},\xi),\xi)^{\top}f(\bar{\theta},\xi)+\frac{L_{\nabla\varphi}L_{f}^{2}}{2}\left\lVert\theta-\bar{\theta}\right\rVert^{2}$
$=\widehat{F}(\theta,\xi;\bar{\theta},I).$
Now we prove Assumption (B4) for Case 1 to Case 3. We prove that the minimum surrogate function $\widehat{F}^{\varepsilon}_{\min}(\cdot,\xi;\bar{\theta})$ satisfies Condition (P4). Assumption (B4) is then verified since $\widehat{F}^{\varepsilon}_{\min}(\cdot,\xi;\bar{\theta})$ by construction is directionally differentiable and satisfies (P2). Since $\widehat{F}^{\varepsilon}_{\min}(\cdot,\xi;\bar{\theta})$ also satisfies the majorizing condition (P2), by the compactness of $\Theta$, it suffices to show that
$\widehat{\kappa}\coloneqq\limsup_{\theta\rightarrow\bar{\theta},\theta\neq\bar{\theta}}\frac{2(\widehat{F}^{\varepsilon}_{\min}(\theta,\xi;\bar{\theta})-F(\theta,\xi))}{\left\lVert\theta-\bar{\theta}\right\rVert^{2}}<+\infty.$
For the piecewise affine function $f(\theta,\xi)\coloneqq g(\theta,\xi)-h(\theta,\xi)$, we have $\mathcal{I}_{h}^{0}(\theta,\xi)\subseteq\mathcal{I}_{h}^{\varepsilon}(\bar{\theta},\xi)$ when $\theta\in\Theta$ is sufficiently close to $\bar{\theta}$. It follows from (P2) that for any $\theta\in\Theta$ sufficiently close to $\bar{\theta}$,
$f(\theta,\xi)\leq\min_{I_{2}\in\mathcal{I}^{\varepsilon}_{h}(\bar{\theta},\xi)}\widehat{f}(\theta,\xi;I_{2})\leq\min_{I_{2}\in\mathcal{I}_{h}^{0}(\theta,\xi)}\widehat{f}(\theta,\xi;I_{2})=f(\theta,\xi).$
Therefore, by similar arguments for lower surrogate functions of $f(\cdot,\xi)$, we obtain that for any $\theta\in\Theta$ sufficiently close to $\bar{\theta}$,
$\min_{I_{2}\in\mathcal{I}^{\varepsilon}_{h}(\bar{\theta},\xi)}\widehat{f}(\theta,\xi;I_{2})=f(\theta,\xi)=\max_{I_{1}\in\mathcal{I}^{\varepsilon}_{g}(\bar{\theta},\xi)}\widecheck{f}(\theta,\xi;I_{1}).$
For Case 1, when $\theta$ is sufficiently close to $\bar{\theta}$, we have
$\widehat{F}^{\varepsilon}_{\min}(\theta,\xi;\bar{\theta})$ $\coloneqq\min_{(I_{1},I_{2},I_{3})\in\mathcal{I}^{\varepsilon}(\bar{\theta},\xi)}\widehat{F}(\theta,\xi;\bar{\theta},I)$
$=\underbrace{\min_{I_{1}\in\mathcal{I}^{\varepsilon}_{g}(\bar{\theta},\xi),I_{2}\in\mathcal{I}^{\varepsilon}_{h}(\bar{\theta},\xi)}\max_{i\in[K_{\varphi}]}\left\{\max\{a_{1i},0\}^{\top}\widehat{f}(\theta,\xi;I_{2})-\max\{-a_{1i},0\}^{\top}\widecheck{f}(\theta,\xi;I_{1})+b_{1i}\right\}}_{\mathcal{V}_{1}}$
$\phantom{==}-\underbrace{\max_{(I_{1},I_{2},I_{3})\in\mathcal{I}^{\varepsilon}(\bar{\theta},\xi)}\left(\max\{a_{2I_{3}},0\}^{\top}\widecheck{f}(\theta,\xi;I_{1})-\max\{-a_{2I_{3}},0\}^{\top}\widehat{f}(\theta,\xi;I_{2})+b_{2I_{3}}\right)}_{\mathcal{V}_{2}}$
$=F(\theta,\xi).$ (24)
The third equality is validated as follows. For the first component $\mathcal{V}_{1}$, we have
$\mathcal{V}_{1}$ $\geq\max_{i\in[K_{\varphi}]}\min_{I_{1}\in\mathcal{I}^{\varepsilon}_{g}(\bar{\theta},\xi),I_{2}\in\mathcal{I}^{\varepsilon}_{h}(\bar{\theta},\xi)}\left\{\max\{a_{1i},0\}^{\top}\widehat{f}(\theta,\xi;I_{2})-\max\{-a_{1i},0\}^{\top}\widecheck{f}(\theta,\xi;I_{1})+b_{1i}\right\}$
$=\max_{i\in[K_{\varphi}]}\left\{\max\{a_{1i},0\}^{\top}\min_{I_{2}\in\mathcal{I}^{\varepsilon}_{h}(\bar{\theta},\xi)}\widehat{f}(\theta,\xi;I_{2})-\max\{-a_{1i},0\}^{\top}\max_{I_{1}\in\mathcal{I}^{\varepsilon}_{g}(\bar{\theta},\xi)}\widecheck{f}(\theta,\xi;I_{1})+b_{1i}\right\}$
$=\max_{i\in[K_{\varphi}]}\left\{a_{1i}^{\top}f(\theta,\xi)+b_{1i}\right\}\eqqcolon\varphi_{1}(f(\theta,\xi),\xi).$
Since there exists $\bar{I}_{1}\in\mathcal{I}^{\varepsilon}_{g}(\bar{\theta},\xi)$ and $\bar{I}_{2}\in\mathcal{I}^{\varepsilon}_{h}(\bar{\theta},\xi)$ such that $\widehat{f}(\theta,\xi;\bar{I}_{2})=f(\theta,\xi)$ and $\widecheck{f}(\theta,\xi;\bar{I}_{1})=f(\theta,\xi)$, we also have
$\mathcal{V}_{1}$ $\leq\max_{i\in[K_{\varphi}]}\left\{\max\{a_{1i},0\}^{\top}\widehat{f}(\theta,\xi;\bar{I}_{2})-\max\{-a_{1i},0\}^{\top}\widecheck{f}(\theta,\xi;\bar{I}_{1})+b_{1i}\right\}$
$=\max_{i\in[K_{\varphi}]}\left\{a_{1i}^{\top}f(\theta,\xi)+b_{1i}\right\}\eqqcolon\varphi_{1}(f(\theta,\xi),\xi),$
indicating $\mathcal{V}_{1}=\varphi_{1}(f(\theta,\xi),\xi)$. Moreover, for the second component $\mathcal{V}_{2}$, when $\theta$ is sufficiently close to $\bar{\theta}$, there exists $\bar{I}_{3}\in\mathcal{I}_{\varphi_{2}}^{\varepsilon}(f(\bar{\theta},\xi),\xi)$ such that $\varphi_{2\bar{I}_{3}}(f(\theta,\xi),\xi)=\varphi_{2}(f(\theta,\xi),\xi)$. Then, we have
$\mathcal{V}_{2}$ $\geq\max\{a_{2\bar{I}_{3}},0\}^{\top}\widecheck{f}(\theta,\xi;\bar{I}_{1})-\max\{-a_{2\bar{I}_{3}},0\}^{\top}\widehat{f}(\theta,\xi;\bar{I}_{2})+b_{2\bar{I}_{3}}=\varphi_{2}(f(\theta,\xi),\xi)$
and
$\mathcal{V}_{2}$ $\leq\max_{(I_{1},I_{2})\in\mathcal{I}_{g}^{\varepsilon}(\bar{\theta},\xi)\times\mathcal{I}_{h}^{\varepsilon}(\bar{\theta},\xi),i\in[K_{\varphi}]}\left(\max\{a_{2i},0\}^{\top}\widecheck{f}(\theta,\xi;I_{1})-\max\{-a_{2i},0\}^{\top}\widehat{f}(\theta,\xi;I_{2})+b_{2i}\right)$
$\leq\max_{(I_{1},I_{2})\in\mathcal{I}_{g}^{\varepsilon}(\bar{\theta},\xi)\times\mathcal{I}_{h}^{\varepsilon}(\bar{\theta},\xi),i\in[K_{\varphi}]}\left\{a_{2i}^{\top}f(\theta,\xi)+b_{2i}\right\}=\varphi_{2}(f(\theta,\xi),\xi),$
indicating $\mathcal{V}_{2}=\varphi_{2}(f(\theta,\xi),\xi)$. Therefore, we derive (7.1) and directly obtain $\widehat{\kappa}=0$. For Case 2, when $\theta$ is sufficiently close to $\bar{\theta}$, we have
$\widehat{F}^{\varepsilon}_{\min}(\theta,\xi;\bar{\theta})$ $\coloneqq\min_{(I_{1},I_{2})\in\mathcal{I}^{\varepsilon}(\bar{\theta},\xi)}\widehat{F}(\theta,\xi;\bar{\theta},(I_{1},I_{2}))$
$=\min_{I_{2}\in\mathcal{I}^{\varepsilon}_{h}(\bar{\theta},\xi)}\varphi^{\uparrow}(\widehat{f}(\theta,\xi;I_{2}),\xi)+\min_{I_{1}\in\mathcal{I}^{\varepsilon}_{g}(\bar{\theta},\xi)}\varphi^{\downarrow}(\widecheck{f}(\theta,\xi;I_{1}),\xi)$
$=\varphi^{\uparrow}\left(\min_{I_{2}\in\mathcal{I}^{\varepsilon}_{h}(\bar{\theta},\xi)}\widehat{f}(\theta,\xi;I_{2}),\xi\right)+\varphi^{\downarrow}\left(\max_{I_{1}\in\mathcal{I}^{\varepsilon}_{g}(\bar{\theta},\xi)}\widecheck{f}(\theta,\xi;I_{1}),\xi\right)$
$=F(\theta,\xi),$
also implying $\widehat{\kappa}=0$. For Case 3, when $\theta$ is sufficiently close to $\bar{\theta}$, we obtain
$\widehat{F}^{\varepsilon}_{\min}(\theta,\xi;\bar{\theta})=\nabla\varphi(f(\bar{\theta},\xi),\xi)^{\top}(f(\theta,\xi)-f(\bar{\theta},\xi))+\varphi(f(\bar{\theta},\xi),\xi)+\frac{1}{2}L_{\nabla\varphi}L_{f}^{2}\left\lVert\theta-\bar{\theta}\right\rVert^{2}.$
Thus, by the second-order Taylor expansion of $\varphi(\cdot,\xi)$ at $f(\bar{\theta},\xi)$, we obtain
$\widehat{F}^{\varepsilon}_{\min}(\theta,\xi;\bar{\theta})-F(\theta,\xi)=\frac{1}{2}L_{\nabla\varphi}L_{f}^{2}\left\lVert\theta-\bar{\theta}\right\rVert^{2}-\frac{1}{2}v^{\top}\nabla^{2}\varphi(f(\bar{\theta},\xi),\xi)v+o(\left\lVert v\right\rVert^{2}),$
@@
8 Details of Convergence Analysis in Section 3.2.3
@@
Proof 8.1
Proof of Proposition 3.4. Given $\varepsilon\geq 0$ and $\rho>0$, for any $\theta^{*}\in\mathcal{S}^{*}$ it holds that $\widehat{F}^{\varepsilon}_{\min}(\theta;\theta^{*})+\frac{\rho}{2}\left\lVert\theta-\theta^{*}\right\rVert^{2}>F(\theta)\geq F(\theta^{*})=\widehat{F}^{\varepsilon}_{\min}(\theta^{*};\theta^{*})$ for any $\theta\neq\theta^{*}$, showing $\theta^{*}\in\mathcal{S}^{d}_{\varepsilon,\rho}$ and thus $\mathcal{S}^{*}\subseteq\mathcal{S}^{d}_{\varepsilon,\rho}$. For any $\theta^{d}\in\mathcal{S}^{d}_{\varepsilon,\rho}$, from Condition (B4) and the first-order optimality condition, it follows that $F^{\prime}(\theta^{d};\theta-\theta^{d})=\widehat{F}^{\varepsilon}_{\min}(\cdot;\theta^{d})^{\prime}(\theta^{d};\theta-\theta^{d})\geq 0$ for all $\theta\in\Theta$, yielding $\theta^{d}\in\mathcal{S}^{d}$ and thus $\mathcal{S}^{d}_{\varepsilon,\rho}\subseteq\mathcal{S}^{d}$.
It remains to show $\mathcal{S}^{d}\subseteq\mathcal{S}^{d}_{0,\rho}$ for any $\rho>0$. Assume by contradiction that there exists $\theta^{d}\in\mathcal{S}^{d}\backslash\mathcal{S}^{d}_{0,\widehat{\rho}}$ for some $\widehat{\rho}>0$. Then there exist $\{\widehat{I}_{s}\in\mathcal{I}^{0}(\theta^{d},\xi^{s})\mid s=1,\dots,n\}$ and $\widehat{\theta}(\neq\theta^{d})\in\Theta$ such that $\widehat{\theta}\in\mathop{\arg\min}_{\theta\in\Theta}\left\{\frac{1}{n}\sum_{s=1}^{n}\widehat{F}(\theta,\xi;\theta^{d},\widehat{I}_{s}\,)+\frac{\widehat{\rho}}{2}\left\lVert\theta-\theta^{d}\right\rVert^{2}\right\}$. By the strong convexity of the function $\theta\mapsto\frac{1}{n}\sum_{s=1}^{n}\widehat{F}(\theta,\xi;\theta^{d},\widehat{I}_{s})+\frac{\widehat{\rho}}{2}\,\left\lVert\theta-\theta^{d}\right\rVert^{2}$, we have $\frac{1}{n}\sum_{s=1}^{n}\widehat{F}(\cdot,\xi;\theta^{d},\widehat{I}_{s})^{\prime}(\theta^{d};\widehat{\theta}-\theta^{d})<0$. Thus, there exists $t_{0}>0$ such that
$F(\theta^{d}+t(\widehat{\theta}-\theta^{d}))\leq\frac{1}{n}\sum_{s=1}^{n}\widehat{F}(\theta^{d}+t(\widehat{\theta}-\theta^{d}),\xi^{s};\theta^{d},\widehat{I}_{s})<\frac{1}{n}\sum_{s=1}^{n}\widehat{F}(\theta^{d},\xi^{s};\theta^{d},\widehat{I}_{s})=F(\theta^{d}),\quad\forall\,t\in(0,t_{0}],$
@@
[…]
@@
Suppose Assumptions (B1)-(B4) hold. For problem (10) with $\max_{\theta\in\Theta}\left\lVert\theta\right\rVert_{\infty}=\mu$, the following statements hold:
@@
    With $\varphi$ in Case 1 or Case 2, $\mathcal{S}^{d}_{\varepsilon,\rho}=\mathcal{S}^{*}$ for any $\varepsilon\in[\varepsilon_{0},+\infty)$ and $\rho\in(0,\frac{M_{\Delta}}{2q\mu^{2}}]$, where $\varepsilon_{0}\coloneqq 2\mu\sqrt{q}\sqrt{\max_{s\in[n]}\left\lVert x^{s}\right\rVert^{2}+1}$, and $M_{\Delta}$ is the constant defined in (EC1).